# Iso-Diffusion: Improving Diffusion Probabilistic Models Using the Isotropy of the Additive Gaussian Noise

### v1

@@ 6.1 Experimental Setup
To validate our approach we consider 2D synthetic data as well as images. For the 2D data, we utilized a conditional dense network consisting of 3 fully-connected hidden layers with ReLU activations. The learning rate was fixed at 1e-3. All the datasets were learned using 1000 time-steps and 1000 epochs.
For the image datasets, we consider the same version of the U-Net utilized in the original DDPM implementation with a learning rate of 2e-4. The U-Net was trained with 1000 time-steps for 1000 epochs as well.
For each dataset, when reporting the metrics, we consider the average of 3 training runs for each of the models. Moreover, all the experiments were run on one Quadro GV-100 GPU with 32GB of VRAM.
6.2 Synthetic Data
For the synthetic generation experiments, we consider four 2D synthetic datasets, namely, the Swiss Roll, the Inter-twining moons, the S-Curve and the 8-gaussians datasets. To evaluate the proposed method, we utilize the most informative generative model metrics, such as the Precision, Recall, Density and Coverage. Of these metrics, Precision and Density measure the fidelity of the generated samples, whereas the Recall and Coverage measure the diversity of the generated samples.
Table 1 summarizes the demonstrated improvements made by our modified loss in terms of these metrics. Across all these datasets we observe that the fidelity metrics, Precision and Density have been improved. Specifically, when considering the Swiss Roll and the Moons datasets, we have been able to achieve much larger improvement in terms of the Density.
Metrics Swiss Roll Moons S-Curve 8-gaussians
DDPM Ours DDPM Ours DDPM Ours DDPM Ours
Precision 0.90 0.982 0.987 0.991 0.989 0.992 0.994 0.997
Recall 0.998 0.984 0.999 0.997 0.996 0.989 0.997 0.988
Density 0.83 0.989 0.957 0.979 0.978 0.982 0.983 0.989
Coverage 0.895 0.917 0.943 0.938 0.938 0.91 0.932 0.90
Table 2: Metrics Variation with the Regularization Parameter for the Swiss Roll Dataset ($\lambda$)
Method Precision Recall Density Coverage
DDPM ([4]) 0.90 (±$\pm 0.005$) 0.998 (±$\pm 0.0001$) 0.83 (±$\pm 0.02$) 0.895 (±$\pm 0.025$)
Ours $\lambda=0.01$ 0.93 (±$\pm 0.022$) 0.998 (±$\pm 0.0002$) 0.885 (±$\pm 0.0035$) 0.90 (±$\pm 0.055$)
Ours $\lambda=0.05$ 0.971 (±$\pm 0.0009$) 0.992 (±$\pm 0.0002$) 0.955 (±$\pm 0.001$) 0.92 (±$\pm 0.015$)
Ours $\lambda=0.10$ 0.982 (±$\pm 0.0001$) 0.984 (±$\pm 0.003$) 0.989 (±$\pm 0.005$) 0.917 (±$\pm 0.001$)
Ours $\lambda=0.20$ 0.99 (±$\pm 0.001$) 0.983 (±$\pm 0.003$) 1.0 (±$\pm 0.001$) 0.875 (±$\pm 0.025$)
Ours $\lambda=0.30$ 0.983(±$\pm 0.003$) 0.96 (±$\pm 0.015$) 1.0 (±$\pm 0.01$) 0.85 (±$\pm 0.035$)
@@ 6.2 Synthetic Data
Table 3: Metrics Variation with the Regularization Parameter for the Moons Dataset ($\lambda$)
Method Precision Recall Density Coverage
DDPM ([4]) 0.987 (±$\pm 0.001$) 0.999 (±$\pm 0.001$) 0.957 (±$\pm 0.009$) 0.943 (±$\pm 0.009$)
Ours $\lambda=0.01$ 0.980 (±$\pm 0.0004$) 0.999 (±$\pm 0.0008$) 0.942 (±$\pm 0.0003$) 0.926 (±$\pm 0.025$)
Ours $\lambda=0.05$ 0.985 (±$\pm 0.004$) 0.998 (±$\pm 0.0001$) 0.957 (±$\pm 0.02$) 0.947 (±$\pm 0.0045$)
Ours $\lambda=0.10$ 0.991 (±$\pm 0.001$) 0.997 (±$\pm 0.001$) 0.979 (±$\pm 0.0001$) 0.938 (±$\pm 0.017$)
Ours $\lambda=0.20$ 0.995 (±$\pm 0.001$) 0.991 (±$\pm 0.0028$) 0.992 (±$\pm 0.005$) 0.924 (±$\pm 0.013$)
Ours $\lambda=0.30$ 0.992 (±$\pm 0.001$) 0.992 (±$\pm 0.005$) 1.0 (±$\pm 0.006$) 0.90 (±$\pm 0.005$)
@@
Table 4: Metrics Variation with the Regularization Parameter for the S-Curve Dataset ($\lambda$)
Method Precision Recall Density Coverage
DDPM ([4]) 0.989 (±$\pm 0.003$) 0.996 (±$\pm 0.003$) 0.978 (±$\pm 0.008$) 0.938 (±$\pm 0.008$)
Ours $\lambda=0.01$ 0.988 (±$\pm 0.003$) 0.996 (±$\pm 0.002$) 0.969 (±$\pm 0.011$) 0.931 (±$\pm 0.01$)
Ours $\lambda=0.05$ 0.989 (±$\pm 0.004$) 0.994 (±$\pm 0.005$) 0.974 (±$\pm 0.011$) 0.925 (±$\pm 0.014$)
Ours $\lambda=0.10$ 0.991 (±$\pm 0.004$) 0.993 (±$\pm 0.005$) 0.977 (±$\pm 0.012$) 0.921 (±$\pm 0.019$)
Ours $\lambda=0.20$ 0.992 (±$\pm 0.004$) 0.989 (±$\pm 0.012$) 0.982 (±$\pm 0.015$) 0.910 (±$\pm 0.036$)
Ours $\lambda=0.30$ 0.993 (±$\pm 0.004$) 0.983 (±$\pm 0.018$) 0.986 (±$\pm 0.017$) 0.898 (±$\pm 0.043$)
Table 5: Metrics Variation with the Regularization Parameter for the 8-gaussians Dataset ($\lambda$)
Method Precision Recall Density Coverage
DDPM ([4]) 0.994 (±$\pm 0.003$) 0.997 (±$\pm 0.001$) 0.983 (±$\pm 0.009$) 0.932 (±$\pm 0.011$)
Ours $\lambda=0.01$ 0.995 (±$\pm 0.002$) 0.997 (±$\pm 0.001$) 0.982 (±$\pm 0.007$) 0.936 (±$\pm 0.011$)
Ours $\lambda=0.05$ 0.995 (±$\pm 0.002$) 0.996 (±$\pm 0.002$) 0.983 (±$\pm 0.006$) 0.931 (±$\pm 0.013$)
Ours $\lambda=0.10$ 0.996 (±$\pm 0.002$) 0.995 (±$\pm 0.003$) 0.986 (±$\pm 0.007$) 0.924 (±$\pm 0.018$)
Ours $\lambda=0.20$ 0.996 (±$\pm 0.002$) 0.992 (±$\pm 0.007$) 0.988 (±$\pm 0.008$) 0.918 (±$\pm 0.022$)
Ours $\lambda=0.30$ 0.997 (±$\pm 0.002$) 0.988 (±$\pm 0.012$) 0.989 (±$\pm 0.009$) 0.90 (±$\pm 0.042$)
(a) DDPM
(b) Ours $\lambda=0.01$
(c) Ours $\lambda=0.05$
(d) Ours $\lambda=0.1$
(e) Ours $\lambda=0.2$
(f) Ours $\lambda=0.3$
Figure 3: Variation of the Generated and Ground Truth Distributions with Changing Regularization Parameter Values
(a) DDPM
(b) Ours $\lambda=0.01$
(c) Ours $\lambda=0.05$
(d) Ours $\lambda=0.1$
(e) Ours $\lambda=0.2$
(f) Ours $\lambda=0.3$
Figure 4: Variation of the Generated and Ground Truth Distributions with Changing Regularization Parameter Values
(a) DDPM
(b) Ours $\lambda=0.01$
(c) Ours $\lambda=0.05$
(d) Ours $\lambda=0.1$
(e) Ours $\lambda=0.2$
(f) Ours $\lambda=0.3$
Figure 5: Variation of the Generated and Ground Truth Distributions with Changing Regularization Parameter Values
The visualized plots of the generated data distributions (figures 1, 3, 4, and 5) indicate that the modified loss function based samples are much more closely gathered and concentrated near the ground-truth data. On the contrary, we can observe that the samples generated via the original DDPM loss function, are not concentrated, but are scattered about. We believe that the isotropy based loss function helps the U-Net learn in such a way that the generated distribution does not contain too many samples which are far away from the most densely packed modes of the ground-truth data distribution. This shows that the proposed loss function, enforces the generated samples to contain properties that pushes them to be closely linked to the real data. Thus, we can directly observe an improvement in the Density metric as it measures the sample fidelity. Furthermore, without any structural information during the training process, the original DDPM is not capable of generating samples which capture the densely packed structure of the distribution and hence it is not necessarily concentrated, but is scattered.
6.3 Unconditional Image Generation
For the unconditional image generation task, we experiment with the Oxford flowers dataset and the Oxford-IIIT Pet dataset. The FID and IS, along with the Precision, Recall, Density and Coverage were used to evaluate the quality of the generated samples.
The results of table 6 further confirm the improvements made by our modified loss on the quality of samples. The Density of the generated images have been significantly improved for the two datasets. Moreover, the FID score has been significanlty improved in the Oxford Flowers dataset by the proposed method. Although the FID and IS are considered to be the most widely used evaluation metrics for assessing image generation, we see that in the case of the Oxford-IIIT-Pet dataset, they convey little to no discerning information about the generative ability of the proposed method and the original DDPM. But, by using other metrics such as the Precision, Recall, Density and Coverage, we can state that while our proposed method suffers a bit in terms of Recall, the generated samples, see figure 2 are very close to being real as indicated by the improvements in the Precision and Density metrics. Moreover, this is further validated by some of the sample generated by both the original DDPM and the proposed method. The samples generated by the proposed method convey much more realistic characteristics in the images than what can be obtained by using the original DDPM. Thus, this should motivate the research community to propose new evaluation metrics such as Density, which is a much more meaningful measure of fidelity over FID and IS, to assess generative models.
Table 6: Comparison of Evaluation Metrics for the two Methods in Unconditional Image Generation
Metrics Oxford Flowers Oxford-IIIT-Pet
DDPM Iso-Diffusion DDPM Iso-Diffusion
FID ($\downarrow$) 49.60 (±$\pm 5.77$) 44.77 (±$\pm 5.77$) 34 (±$\pm 0.76$) 34.47 (±$\pm 0.17$)
IS ($\uparrow$) 3.93 (±$\pm 0.12$) 3.94 (±$\pm 0.12$) 12.0 (±$\pm 0.27$) 12.0 (±$\pm 0.27$)
Precision ($\uparrow$) 0.75 (±$\pm 0.01$) 0.86 (±$\pm 0.01$) 0.71 (±$\pm 0.03$) 0.80 (±$\pm 0.03$)
Recall ($\uparrow$) 0.21 (±$\pm 0.06$) 0.14 (±$\pm 0.01$) 0.29 (±$\pm 0.02$) 0.24 (±$\pm 0.22$)
Density ($\uparrow$) 2.48 (±$\pm 0.99$) 3.99 (±$\pm 0.35$) 2.11 (±$\pm 0.22$) 2.99 (±$\pm 0.37$)
Coverage ($\uparrow$) 0.93 (±$\pm 0.06$) 0.97 (±$\pm 0.01$) 0.98 (±$\pm 0.01$) 0.99 (±$\pm 0.01$)
6.4 Variation of the Metrics with the Regularization Parameter
Although the performance of the modified loss function has been able to produce sample which surpass the original DDPM’s samples quality, the quality depends on the regularization parameter of the modified loss function. In particular, we performed a few more experiments by considering a range of values for the regularization parameter.
@@ 6.4 Variation of the Metrics with the Regularization Parameter
The metrics for the 2D synthetic datasets with different values of the regularization parameter ranging from 0.01 to 0.30 are tabulated in tables 2, 3, 4, and 5. We see that the fidelity metrics, Precision and Density, gradually improve with the increase of the regularization parameter. On the other hand, however, we can see that the diversity metrics, Recall and Coverage, gradually decline with the parameter. We believe that this is a direct consequence of imposing a structural constraint on the objective function. It is evident that by focusing on the structure or the isotropy of the distribution, our method is capable of capturing highly dense mode regions and generating samples near them rather than being too diverse. Thus, it increase the fidelity but decreases the diversity of the generated samples.
@@ 7 Conclusion
Denoising Diffusion Probabilistic Models have achieved state-of-the-art performance in generative modeling tasks such as unconditional image generation and image super resolution. However, these models can still be improved upon and much work has been put into improving them. In this paper, we propose another improvement method which is built on the premise that since the distribution that the forward process terminates and the reverse process initiates at a white gaussian distribution, which is isotropic and is void of any structure, it is well motivated, that the incorporation of isotropy as a measure of structure on the loss function will improve the DDPMs’ generated sample fidelity. We, theoretically, show that isotropy is well a defined metric to measure the structure of the distribution during the forward process and the proposed modification helps the DDPM to converge to better solutions based on the simple modified loss. Finally, we validate and show that our modified objective function improves the performance of the DDPM, via experiments performed on 2D synthetic datasets and on unconditional image generation.

### v2

@@ 7.1 Experimental Setup
To validate our approach we consider 2D synthetic data as well as images. For the 2D data, we utilized a conditional dense network consisting of 3 fully-connected hidden layers with ReLU activations. The learning rate was fixed at 1e-3. All the datasets were learned using 1000 time-steps and 1000 epochs. For the image datasets, we consider the same version of the U-Net utilized in the original DDPM implementation with a learning rate of 2e-4. The U-Net training involved 1000 time-steps and 1000 epochs for the Oxford Flower and Oxford-IIIT Pet datasets, while the CIFAR-10 and CIFAR-100 datasets were trained with 2000 epochs. Metrics are reported as the average of 3 training runs per dataset, with PRDC values calculated using k=5 nearest neighbors for each dataset. Moreover, all the experiments were run on one Quadro GV-100 GPU with 32GB of VRAM.
7.2 Performance comparison between DDPM and Iso-Diffusion
To compare the performance between DDPM and Iso-Diffusion, the modified loss function was utilized in all four 2D synthetic datasets, Oxford Flower dataset, Oxford IIIT Pet dataset and CIFAR-10 dataset. Precision, Recall, Density along with Coverage were used to evaluate and compare the performance of the two models on 2D synthetic datasets. In addition to those four evaluation metrics, FID and IS were used to evaluate the quality of the generated samples for the image datasets Oxford Flower, Oxford-IIIT Pet, CIFAR-10 and CIFAR-100.
Table 1 demonstrates the comparison between the best performing isotropy based model (Iso-Diffusion) and DDPM in terms of the generative model’s evaluation metrics along with the percentage change from DDPM. Across all these 2D synthetic datasets we observed that the fidelity metrics, Precision and Density have been improved in Iso-Diffusion. The results of Table 2 further confirm the improvements made by our modified loss on the quality of image samples. The Density of the generated images has been significantly improved for all four datasets. Moreover, the FID score has been significantly improved in the CIFAR-10 dataset by the proposed method.
Although the performance of the modified loss function has been able to produce sample which surpass the original DDPM’s samples quality, the quality depends on the regularization parameter of the modified loss function. In particular, we performed a few more experiments by considering a range of values for the regularization parameter. The metrics for the Oxford Flower dataset and Oxford-IIIT-Pet dataset with different values of the regularization parameter ranging from 0.01 to 0.30 are tabulated in Table 3 and Table 4. We see that the fidelity metrics, Precision and Density, gradually improve with the increase of the regularization parameter. However, we can see that the diversity metrics, Recall and Coverage, gradually decline with the parameter.
Although the FID and IS are considered to be the most widely used evaluation metrics for assessing image generation, we see that in the case of all four datasets, they convey little to no discerning information about the generative ability of the proposed method and the original DDPM. But, by using other metrics such as the Precision, Recall, Density and Coverage, we can state that while our proposed method suffers a bit in terms of Recall, the generated samples, are very close to being real (see Figure 1), as indicated by the improvements in the Precision and Density metrics.
7.3 Interpretation of the results of 2D data distributions using PRDC values
Method FID ($\downarrow$) IS ($\uparrow$) Precision ($\uparrow$) Recall ($\uparrow$) Density ($\uparrow$) Coverage ($\uparrow$)
DDPM 55.5900 3.0970 0.7248 0.1840 2.6320 0.9588
Ours $\lambda$ = 0.01 53.3374 3.2023 0.7839 0.1570 3.3445 0.9758
Ours $\lambda$ = 0.05 54.7064 3.2208 0.733 0.17625 2.6384 0.9586
Ours $\lambda$ = 0.10 47.3097 3.5037 0.9441 0.0555 11.0389 0.9935
Ours $\lambda$ = 0.30 51.5820 3.3105 0.9460 0.0549 12.5441 0.9946
@@ 7.3 Interpretation of the results of 2D data distributions using PRDC values
Table 3: Metrics Variation with the Regularization Parameter for the Oxford Flower Dataset ($\lambda$)
Method FID ($\downarrow$) IS ($\uparrow$) Precision ($\uparrow$) Recall ($\uparrow$) Density ($\uparrow$) Coverage ($\uparrow$)
DDPM 34.0874 7.0827 0.8189 0.1522 6.7040 0.9996
Ours $\lambda$ = 0.01 32.7278 7.5298 0.8805 0.1233 7.9743 0.9991
Ours $\lambda$ = 0.05 32.4877 7.5156 0.8628 0.1263 8.0348 0.9997
Ours $\lambda$ = 0.10 33.3405 7.4813 0.9103 0.1015 10.6351 1.0000
Ours $\lambda$ = 0.30 31.8998 7.5313 0.9542 0.0633 15.7780 0.9999
@@
Table 4: Metrics Variation with the Regularization Parameter for the Oxford-IIIT-Pet Dataset ($\lambda$)
Figure 4: Center Banana 2D synthetic dataset. (a) True distribution points, color-coded by k=5 nearest neighbor radius. (b) DDPM-generated points, color-coded by true manifold span per point. (c) Iso-Diffusion generated points, color-coded by true manifold span per point.
Figure 5: Scattered Moon 2D synthetic dataset. (a) True distribution points, color-coded by k=5 nearest neighbor radius. (b) DDPM-generated points, color-coded by true manifold span per point. (c) Iso-Diffusion generated points, color-coded by true manifold span per point.
Figure 6: Swiss Roll 2D synthetic dataset. (a) True distribution points, color-coded by k=5 nearest neighbor radius. (b) DDPM-generated points, color-coded by true manifold span per point. (c) Iso-Diffusion generated points, color-coded by true manifold span per point.
@@
We believe that the disparity in the changes of Precision, Recall, Density and Coverage is a direct consequence of imposing a structural constraint on the objective function. It is evident that by focusing on the structure or the isotropy of the distribution, our method is capable of capturing highly dense mode regions and generating samples near them rather than being too diverse. Thus, it increases the fidelity but decreases the diversity of the generated samples.
As illustrated in the Figure 4(a), the Central Banana distribution was designed by introducing a distinct mode to the main structure of the distribution resulting in a multimodal distribution with a density gradient. Once, it is generated via Iso-Diffusion as indicated in Figure 4(c), it is evident that, Iso-Diffusion, is capable of capturing the main structure even with the discontinuities of the density gradient. However, the illustrations show that DDPM lacks the capability of capturing the discontinuity in the density gradient between the tail end of the main distribution and the distinct mode. Instead, it tries to generate data points that are actually not even in the true distribution by interpolating along the main lobe’s trend (see Figure 4(b)). Moreover, the limited capability to capture the discontinuity in the density gradient of DDPM can be further observed in the Swiss Roll distribution as well (see Figure 6(b) and 6(c)). The increase in Density and decrease in Coverage for the datasets Swiss Roll and Central Banana are clear evidence for the aforementioned observations. Hence, it is limited in ability to capture the underlying structure of the distribution. Additionally, there is a noticeable trend of generating data points (blue points in Figure 4(b), 4(c)), outside the boundaries of the highly dense regions of the main lobe. This effect is likely due to the model’s focus on these high-density regions. However, compared to DDPM, Iso-Diffusion effectively regulates the overgeneration of data points outside the boundaries of densely packed regions. This improvement is likely a result of the added regularization in the improved object function, which encourages capturing the main semantics of the true distribution.
As illustrated in the Figure 5(a), the Scattered Moon distribution was designed by imposing scattered noise around the main structure of the data distribution. Once, it is generated via Iso-Diffusion as indicated in the Figure 5(c), it is evident that, the model has tried to only capture the underlying semantics of the distribution without being susceptible to the low probable regions. Whereas, the DDPM model shows limitations in capturing the distinction between the main structure and the scattered noise (see Figure 5(b)). The increased Density and reduced Coverage values support this observation. This shows that the proposed objective function, enforces the generated samples to contain properties that push them to be closely linked to the real data. Thus, we can directly observe an improvement in the Density metric as it measures the sample fidelity. We believe that in the context of unconditional image generation, the isotropy based objective function helps the U-Net learn to keep the generated samples closer to the high-density regions of the ground-truth distribution.
These observations highlight the proposed algorithm’s ability to increase Density by focusing on the dense regions of the true distribution. At the same time, the absence of generated data in the neighborhoods of low probable data points in the true distribution may result in a reduction in Coverage. When scattering is minimal, Coverage remains consistent. This indicates that the algorithm effectively captures the main structure of the true distribution without extending into low probable regions. Also, each of these metrics has their own utility depending on the application [13]. Thus, this should motivate the research community to propose new evaluation metrics such as Density, which is a much more meaningful measure of fidelity over FID and IS, to assess generative models.
Preserving the modality of a data distribution is essential, as failing to capture it can lead to a loss of semantic details or edge information, both of which represent high-level features in computer vision and image processing tasks [4], [29]
@@ 8 Conclusion
Denoising Diffusion Probabilistic Models have achieved state-of-the-art performance in generative modeling tasks such as unconditional image generation and image super-resolution. However, these models can still be improved upon and much work has been put into improving them. In this paper, we propose another improvement method that is built on the premise that since the distribution that the forward process terminates and the reverse process initiates at an isotropic Gaussian distribution, which is void of any structure, it is well motivated, that the incorporation of isotropy as a measure of structure on the loss function will improve the DDPMs’ generated sample fidelity. We, theoretically, show that isotropy is well a defined metric to measure the structure of the distribution during the forward process and the proposed modification helps the DDPM to converge to better solutions based on the simple modified loss. We show that the Iso-Diffusion objective function regulates data point generation, aligning it with the prominent structures of the true distribution and anchoring the process to dense, information-rich modes. Finally, we validate and show that our modified objective function improves DDPM performance through experiments on 2D synthetic datasets and unconditional image generation, supported by an in-depth analysis and evidenced by improved fidelity and diversity metrics.